Nov 3, 13:00 UTC Fully fixed - We have implemented a fix for both the DNS and switching problems, all systems are fully operational. Nov 3, 07:23 UTC Partial Fix - The problems with two switches has been solved. One of the DNS clusters is online again, so that nameserver resolution should be working. Nov 3, 06:59 UTC Identified - According to our NOC two switches are affected, also. No ETA yet. Nov 3, 06:38 UTC Identified - We are observing a partial outage of our infrastructure due to a dDoS targeting our nameservers.